Operations Report – January 13th



Crew 272 Operations Report 13-1-2023

SOL: 12

Name of person filing report: Mason Kuhn

Non-nominal systems: Oven light

Notes on non-nominal systems: No additional issues


Spirit rover used: Yes

Hours: 207.3

Beginning charge: 100%

Ending charge: 73%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.0

Beginning charge: 100%

Ending charge: 78%

Currently charging: Yes

Curiosity rover used: No

Hours: 219.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Spirit and Opportunity were used on the EVA. They performed nominally. The chargers were draped over the front bumper upon return to the Hab. Perseverance is still at the outpost temporarily.

Summary of Hab operations: Post-sim cleanings to be conducted in the afternoon (upper deck, RAM, Sci Dome, GreenHab), and cleanings of the lower deck will be completed this afternoon with touch-up cleanings taking place in the morning of Sol 13.

WATER USE: 11.8 gallons (by 2:30pm and before cleanings that use water)

Water (static tank): 142.8 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suits 1, 5, 9, and 10 were used on the EVA. All suits performed nominally, and all are cleaned & currently charging. All radios have been charged to full, and they have been removed from the charging docks. The EVA crew had no communication issues.

Summary of GreenHab operations: Monitored and watered GreenHab crops, with a bit of extra watering due to higher temperatures. Fertilized in the afternoon. Harvested beans, chives, thyme, and dill. Planted cilantro and radishes. Cleared out dead leaves and overgrowth, distributed organic matter around GreenHab to recycle nutrients, and swept and cleaned the floor & cabinets.

WATER USE: 9 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 15g beans, 3g chives, 15g thyme, 30g dill

Summary of Science Dome operations: Crew items were removed from the Science Dome, and cleanings according to the checklist to be completed shortly

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: All tools from the RAM that were in the lower deck have been returned to the RAM after final measurements were made, cleanings were conducted.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Copyright © The Mars Society. All rights reserved. | Main Site